OpenStudy (anonymous):

simplify the expression square root -16/(3-3i)+(1-2i)

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{-16}}{(3-3i)+(1-2i)}?\)

OpenStudy (anonymous):

CORRECT

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{-16}}{(3-3i)+(1-2i)}\implies \cfrac{\sqrt{-1\cdot 4^2}}{3-3i+1-2i} \\ \quad \\ \cfrac{\sqrt{-1}\cdot \sqrt{4^2}}{4-5i}\implies \cfrac{ 4i}{4-5i}\) and surely you've done this ones already.... get the conjugate and multiply top and bottom by that to get rid of the "i" at the bottom

OpenStudy (anonymous):

ok would it be 8-4i/15?

OpenStudy (jdoe0001):

well.. not quite.... did you find the conjugate of the denominator?

OpenStudy (anonymous):

is it close or am i way off

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

well... hmmm how did you get it though?

OpenStudy (jdoe0001):

any ideas on the conjugate of the denominator?

OpenStudy (anonymous):

nope. can you tech me how to find the conjugate of the denominaotr

OpenStudy (jdoe0001):

hmmm did you follow the above? confused about it yet?

OpenStudy (anonymous):

yeah i understand all of the above. but there has to be another step

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/images/conjugate.gif <--- see what the conjugate is? what do you think it'd be for say the denominator in this case?

OpenStudy (anonymous):

so the sign just flips?

OpenStudy (anonymous):

so the comjugate would be 4+5i?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

ok thank you i got it

OpenStudy (jdoe0001):

thus \(\bf \cfrac{ 4i}{4-5i}\cdot \cfrac{4+5i}{4+5i}\implies \cfrac{ 4i({4+5i})}{(4-5i)({4+5i})} \\ \quad \\ recall\implies {\color{brown}{ (a-b)(a+b) = a^2-b^2}}\qquad thus \\ \quad \\ \cfrac{ 4i({4+5i})}{(4-5i)({4+5i})}\implies \cfrac{ 4i({4+5i})}{(4)^2-(5i)^2}\implies \cfrac{ 4i({4+5i})}{16-5^2i^2}\) can you see it from there?

OpenStudy (anonymous):

-20-16i/9

OpenStudy (jdoe0001):

well.... close.. lemme finish it

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf \cfrac{ 4i}{4-5i}\cdot \cfrac{4+5i}{4+5i}\implies \cfrac{ 4i({4+5i})}{(4-5i)({4+5i})} \\ \quad \\ recall\implies {\color{brown}{ (a-b)(a+b) = a^2-b^2}}\qquad thus \\ \quad \\ \cfrac{ 4i({4+5i})}{(4-5i)({4+5i})}\implies \cfrac{ 4i({4+5i})}{(4)^2-(5i)^2}\implies \cfrac{ 4i({4+5i})}{16-5^2i^2} \\ \quad \\ \cfrac{16i+20i^2}{16-25i^2} \\ \quad \\ recall\implies {\color{brown}{ i^2\to -1}}\qquad thus \\ \quad \\ \cfrac{16i+20i^2}{16-25i^2}\implies \cfrac{16i+20(-1)}{16-25(-1)}\implies \cfrac{-20+16i}{16+25} \\ \quad \\ \cfrac{-20+16i}{41}\)